# On denominators of consecutive $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions

### v1

@@ 5. An index identity and the $\Phi$ function
As $a_{i+1}q_{i}-a_{i}q_{i+1}=1$, we have
$\nu(\gamma_{i}):=\frac{q_{i-1}+q_{i+1}}{q_{i}}=\frac{a_{i-1}+a_{i+1}}{a_{i}}=a_{i+2}q_{i}-a_{i}q_{i+2}\in\mathbb{N},\qquad i=1,\ldots,S_{Q}.$
The proof of Proposition 3 follows closely the idea of [10, Theorem 1] and proceeds by induction on $Q$. Denote the sum from (4) by $T_{Q}$. For $Q=3$, we have $\mathscr{S}_{3}=\{1/1\}$, $S_{3}=1$, and $T_{3}=2=3R_{3}-1$. Suppose that $T_{Q-1}=3S_{Q-1}-1$ for some $Q\geq 4$. The number of mediant insertions when enlarging $\mathscr{S}_{Q-1}$ to $\mathscr{S}_{Q}$ is precisely $S_{Q}-S_{Q-1}$. When such a mediant insertion $\gamma_{\ast}=a_{\ast}/q_{\ast}$ occurs, say between $\gamma_{i}$ and $\gamma_{i+1}$, its contribution to $T_{Q-1}$ is
@@
Figure 4. The horizontal and vertical transformations when insertions are made. Thus, a point $(q_{1},q_{2})\in\mathcal{D}_{Q}$ that disappears at level $Q$ is replaced in $\mathcal{D}_{Q+1}$ with two new points $(q_{1}+q_{2},q_{2})$ and $(q_{1},q_{1}+q_{2})$. Therefore, through insertions, the polygon $ABEF$, in which the points that will disappear at level $Q$ are found, transforms into the polygons $ABCD$ and $GDEF$ that will contain the new points.
@@
In Figures 3, 4, and 5, it can be observed that not only does $\Phi(Q)$ exhibit significant oscillation, but also that at times, particularly around levels $Q$ with a large number of divisors, the disappearing and newly created points tend to group together into distinct formations.

### v2

@@ 5. An index identity and the $\Phi$ function
The proof of Proposition 3 follows closely the idea of [11, Theorem 1] and proceeds by induction on $Q$. As $a_{i+1}q_{i}-a_{i}q_{i+1}=1$, we have
$\frac{q_{i-1}+q_{i+1}}{q_{i}}=\frac{a_{i-1}+a_{i+1}}{a_{i}}=a_{i+2}q_{i}-a_{i}q_{i+2}\in\mathbb{N},\qquad i=1,\ldots,S_{Q}.$
Denote
$T_{Q}:=\sum\limits_{i=1}^{S_{Q}}\frac{q_{i-1}+q_{i+1}}{q_{i}}.$
For $Q=3$, we have $\mathscr{S}_{3}=\{1/1\}$, $S_{3}=1$, and $T_{3}=2=3S_{3}-1$. Suppose that $T_{Q-1}=3S_{Q-1}-1$ for some $Q\geq 4$. The number of mediant insertions when enlarging $\mathscr{S}_{Q-1}$ to $\mathscr{S}_{Q}$ is precisely $S_{Q}-S_{Q-1}$. When such a mediant insertion $\gamma_{\ast}=a_{\ast}/q_{\ast}$ occurs, say between $\gamma_{i}$ and $\gamma_{i+1}$, its contribution to $T_{Q-1}$ is
@@
Figure 4. The horizontal and vertical transformations when insertions are made. Thus, a point $(q_{1},q_{2})\in\mathscr{D}_{Q}$ that disappears at level $Q$ is replaced in $\mathscr{D}_{Q+1}$ with two new points $(q_{1}+q_{2},q_{2})$ and $(q_{1},q_{1}+q_{2})$. Therefore, through insertions, the polygon $ABEF$, in which the points that will disappear at level $Q$ are found, transforms into the polygons $ABCD$ and $GDEF$ that will contain the new points.
@@
In Figures 34, and 5, it can be observed that not only does $\Phi(Q)$ exhibit significant oscillation, but also that at times, particularly around levels $Q$ with a large number of divisors, the disappearing and newly created points tend to group together into distinct formations.